BG February Com Night 2, UT February 2 2009

Observers: D.Thompson (in Tucson, during Super Bowl), and then J. Hill and A. Rakich (in Tucson, after 9PM), and N. Ageorges (in Garching, around midnight) and J. Storm (in Potsdam, after 1AM)
Telescope Operator: A. Ceranski
SW Support: M. Delapena
Telescope Support:

Summary:

Quite a calm night. Build 18 seems to work as expected, including LSS.

Seeing was horrible in the middle of the night. Improving to 0.6 arcsec for the last 2 hours.

Main telescope problem was SX Primary failing to meet tolerance about 15% of the time. It looks like a problem with HP3.

Still have AGW transformation issues. Took a new set of calibration data for this.

Also took some fast WFS data.

Preparations:

Butch and friends have revived the winter chiller, so we are able to cool the oil and run left mirror ventilation.

Details (Times in MST):

17:06 Both check4nport and oacserver found running on agw-control. Homed motors without problems.

17:25 Restarting TCS after switch to B18.

17:42 Encoders completed resetting.

17:48 Checking 107 to see if GCSL is the only thing running there. Found the following, what is it?
root      3126     1  0  2008 ?        00:00:10 sendmail: accepting connections
smmsp     3134     1  0  2008 ?        00:00:00 sendmail: Queue runner@01:00:00 for /var/spool/clientmqueue

and LSS, GCSL, and PMCL all running on 107. Will stop and restart all of TCS to fix ownership issues and make sure only GCS is running on 107. LSS started normally, and ran the remainder of the night.

18:14 irc Authorize 'bentGregorianFront left' LUCIFER
The authorize was coming back OK even though LBC was authorized on the left side. The same command to both fails as expected.

18:16 Open Dome (Left Shutter Door only), Tout=-3.4 degC. Dewpt=-11.3 degC, RH=55% Wind (LBTO) 10-15 m/sec. M1=-0.9C. Running on winter chiller. Side Vent Doors open, and Rear Vent Doors open midway.

Verification of B18 Active Optics Modes

18:17 pointto WT10_227 Tel="ACQUIRE" GS=0 pam-. IE=-20 CA=-12 (about 25 arcsec offset)

18:31 pointto BS9105 ACTIVE on axis, succeeded. Collimation sent to M1. Check mode 3d and old (pure M1) performance.
  • Guider is running 2s exposures, 0.15 effecitve Hz guide rate
  • Standard (old) mode: All active optics sent to M1 (guiderimage through 3302)
  • Sign test: Z4 offload to M2 in PCS mode 3d - focus is stable
  • Mode 3d: Z11 and Z22 applied with focus portion applied to M2. (~3310 through 3562)
    M1 total colimation: X=-0.058 Y=-0.234 Z=0.247 RX=-13.043 RY=-4.469 M1Temp=-1.53
  • Useful to repeat 3b (Z11, Z22 FCIN=1) for comparison (guiderimage 3588 through ~3947)
    Collimate button clicked at around 3707 to clear scrolling errors on PSF but things were OK.
  • End of tests at EL~50.

20:24 pointto BS9126 ACTIVE on axis

  • Back in 3d mode. Suspend guiding offsets from being applied (TRACK mode). WFS stopped.
    (guiderimage 3965 through 4200)

Torsten: it looks like the auto-display of acquisition images is only displaying 512^2 pixels (or less) and thus not displaying the full active area od the CCD since we could see the left overscan but not the bottom stripe, and we could not see the star at the edge of the image in acquire000025.fits. Also, the title on the displayed acquisition image was #6, not #25.

20:37 sent stepfocus 0.0 to clear off secondary offload. Suspend application of guiding offsets (guiderimage 4106- )
hit collimate button at: 20:45:10, Suspend WFS at 20:41 (#4106).

WinCC reads HBS oil temp of ((9.43-10.) * 10.) = -5.7C. Tout=-4.3 TM1=-2.9

Pittsburgh Steelers 27 -- Arizona Cardinals 23

21:00 A happy JMH takes over after a very close Super Bowl game decided in the last minute of the game!

21:10 We had been guiding, but without WFS we were drifting out of focus.

21:11 pointto BS9126 ACTIVE on axis - That "improves" the seeing to 1.6 - 2.0 arcsec.

21:15 Michele and Dave T. are cleaning out disk space on /var on lbtmu110. This may or may not have been related to the LSS problems last night.

21:20 CH0401 with HBS, IC and SX mirror ventilation is holding -16.57 degC LWT and -12.93 degC EWT.

Fast WFS Data

Seeing is varying between 2 and 3.5 arcsec, so it isn't clear that this data is worth much. "like trying to do landscape painting in fog" (Rakich). Sec Act Opt Offload is Yes (but we haven't offloaded in half an hour), and Z11/Z22 focus compensation is on M2 (and we haven't updated M1 collimation since EL=66.7 deg, now EL=79 deg)

21:51 Stop WFS so Andrew can collimate in IDL through wfscimage 000006

Directory 090202w_fast

21:54 Fast Data wfscimage 000007 - 18, Twelve images of 5 sec exp

21:58 Collimating wfscimage 19 - 25

22:03 Fast Data wfscimage 000026 - 37, Twelve images of 5 sec exp

22:07 Collimating wfscimage

22:09 Increase collimation exposure time from 20 to 30 sec.

22:21 Fast Data wfscimage 000052 - 63, Twelve images of 5 sec exp

22:2x Collimating

22:32 Fast Data wfscimage 000071 - 82, Twelve images of 5 sec exp

22:41 Fast Data wfscimage 000089 - 100, Twelve images of 5 sec exp

22:45 5 Collimation loops with IDL, 30 sec exposures.

22:50 guiding through transit, now above 86 degEL

22:56 Restart GCS WFS

Exercising offsets from IRAF

23:12 offset 5 0 coord="DETXY" --> offset_x=3, x=3

23:16 offset 5 0 coord="DETXY" (verifying the default ABS)

23:20 focus 0 - M1 chase failed to make tolerance, but pressing collimate fixes it almost instantly.

23:24 offset 5 0 type="REL" coord="DETXY" --> offset_x=6, x=6

23:26 offset 5 2 type="REL" coord="DETXY"

(Note that the lpar labels in LBTtools.IRTC.offset imply millimeters for DETXY offsets, while Jose is processing all values as arcsec regardless of the coordinate system.)

23:28 Aaron reports 13 arcsec seeing on Polaris monitor, while we are seeing 2-3 arcsec on the guider.

00:00 Block guiding corrections from PCS. GCSGUI issues a nice warning. The image still shows 1 arcsec atmospheric excursions from one guide frame to the next. This would be a cause where you might want to lower the gain on guide corrections to average over several minutes.

00:06 Resume PCS Guiding

00:07 Absorb DETXY offsets from PCS GUI.

00:27 Seeing on the guider continues in the range of 2-3.5 arcsec.

00:32 focus 0

01:00 Now that Jesper has joined us, the Polycom topology is even more complex. We (LBTO Software room ...46) and Jesper are connected to the LaPlace unit in 505A (128.196.208.238); Aaron and Nancy are connected to the N505 Conference Room (128.196.208.27), and the two conference rooms are connected together.

01:17 focus 0

LUCIFER tests

01:24 preset ACTIVE with LUCIFER -succeeded.

Offsets to Test Transformations

01:42 starting the offset script from LUCIFER

Offset command took 33 sec, and stepFocus command took 38 sec.
Mon Feb  2 01:43:41.501 2009  Offset command started for left LUCIFER from IIF
Mon Feb  2 01:44:15.264 2009  Offset command complete for left LUCIFER from IIF
Mon Feb  2 01:44:15.311 2009  StepFocus command started for left LUCIFER for 0.00000
Mon Feb  2 01:44:53.909 2009  StepFocus command complete for left LUCIFER for 0.00000

01:56 preset ACTIVE with LUCIFER succeeded

02:14 Reinstalled the correct LUCIFER focus value of 31.0 in AIP_L.cfg. Why was the wrong value distributed with Build 18?

Mon Feb 2 02:15:29.579 2009 StepFocus alert for left LUCIFER: left PMC Mirror GoToPositionAbs Command Failed because Chase failed to make tolerance.
Why is the SX Primary having this problem?

02:17 Preset ACTIVE with LUCIFER succeeded

Guider is 1.7 arcsec, but LUCIFER is getting 1.3 arcsec in K.

02:32 Preset ACTIVE with LUCIFER succeeded. PA=-50

02:51 Preset ACTIVE with LUCIFER succeeded.

Mon Feb 2 03:08:52.405 2009 Preset alert for left LUCIFER mode ACTIVE: left PMC Mirror GoToPositionAbs Command Failed because Chase failed to make tolerance

03:12 Seeing has improved to 1.2 arcsec on the guider.

Mon Feb 2 03:25:19.740 2009 Preset alert for left LUCIFER mode ACTIVE: left PMC Mirror GoToPositionAbs Command Failed because Chase failed to make tolerance

03:33 Only 5 minutes after Andrew leaves the building, the seeing has improved to 1 arcsec (briefly).

Mon Feb 2 03:49:15.137 2009 Hardpoint 3 lvdt is not showing any motion
Mon Feb 2 03:49:16.422 2009 Hardpoint 3 lvdt is not showing any motion
Mon Feb 2 03:49:17.453 2009 Mirror GoToPositionAbs Command Failed because Chase failed to make tolerance

04:00 Adjust IE=-20.0 CA=-19.55

Mon Feb 2 04:01:05.994 2009 Preset alert for left LUCIFER mode TRACK: left PMC Mirror GoToPositionAbs Command Failed because Chase failed to make tolerance

PCS to AGW Calibration

This is real transcontinental observing. John is presetting the telescope. Jesper is recording the probe coordinates. Nancy is taking the LUCIFER images.

04:17 pointto BS9152 gs=2 PA=0 TelM="ACTIVE" (acquire 36)

04:25 pointto BS9152 gs=4 PA=0 TelM="ACTIVE" (acquire 37)

04:28 pointto BS9152 gs=5 PA=0 TelM="ACTIVE" - out of range (there are two star #5s on the chart)

04:30 pointto BS9152 gs=7 PA=0 TelM="ACTIVE" (acquire 38)

04:35 pointto BS9152 gs=8 PA=0 TelM="ACTIVE" (acquire 39) (binary)

04:39 pointto BS9152 gs=10 PA=0 TelM="ACTIVE" (acquire 40)

04:41 pointto BS9152 gs=14 PA=0 TelM="ACTIVE" (acquire 41)

04:44 pointto BS9152 gs=15 PA=0 TelM="ACTIVE" (acquire 42)

04:47 pointto BS9152 gs=18 PA=0 TelM="ACTIVE" (acquire 43) - failed to get star in guiding box

04:49 pointto BS9152 gs=19 PA=0 TelM="ACTIVE" - out of range

Seeing has improved to 0.6 arcsec on guider.

04:50 pointto BS9152 gs=2 PA=0 TelM="ACTIVE" (acquire 44)

04:55 pointto BS9152 gs=4 PA=0 TelM="ACTIVE" (acquire 45)

04:58 pointto BS9152 gs=14 PA=0 TelM="ACTIVE" (acquire 46)



05:03 pointto BS9153 gs=2 PA=0 TelM="ACTIVE" - out of range?

05:05 pointto BS9153 gs=3 PA=0 TelM="ACTIVE" - probe wouldn't move

05:07 pointto BS9153 gs=14 PA=0 TelM="ACTIVE" - probe wouldn't move

05:07 Home the probe

05:09 pointto BS9153 gs=14 PA=0 TelM="ACTIVE" (acquire 47) - Stray light on the WFS, so the images are ugly.



05:19 pointto BS9145 gs=20PA=0 TelM="ACTIVE" - couldn't move probe

05:19 pointto BS9145 gs=18 PA=0 TelM="ACTIVE" - couldn't move probe

05:20 Home the probe

05:22 pointto BS9145 gs=18 PA=0 TelM="ACTIVE" - couldn't move probe

05:23 pointto BS9145 gs=0 PA=0 TelM="ACTIVE" - couldn't move probe

05:24 Home the probe

05:29 stopAGW, startAGW, Home the probe

05:34 pointto BS9145 gs=0 PA=0 TelM="ACTIVE" - works OK but no data on-axis

05:37 pointto BS9145 gs=18 PA=0 TelM="ACTIVE" - no star found (much too faint)

05:39 pointto BS9145 gs=16 PA=0 TelM="ACTIVE" (acquire 50)

05:44 pointto BS9145 gs=8 PA=0 TelM="ACTIVE" (acquire 51)

05:47 pointto BS9145 gs=6 PA=0 TelM="ACTIVE" (acquire 52)

05:50 pointto BS9145 gs=9 PA=0 TelM="ACTIVE" (acquire 53)

05:53 pointto BS9145 gs=4 PA=0 TelM="ACTIVE" (acquire 54) - no star found (too faint)

05:57 pointto BS9145 gs=14 PA=180 TelM="ACTIVE" (acquire 55) - no star found (too faint)

05:58 pointto BS9152 gs=12 PA=180 TelM="ACTIVE" (acquire 56)

Mon Feb 2 05:59:30.683 2009 Preset alert for left LUCIFER mode ACTIVE: left PMC Mirror GoToPositionAbs Command Failed because Chase failed to make tolerance Mon Feb 2 05:59:30.701 2009 Preset alert for left LUCIFER mode ACTIVE: Ok

06:04 pointto BS9152 gs=17 PA=180 TelM="ACTIVE" -acquisition offset failed (acquire 57)

06:06 pointto BS9152 gs=17 PA=180 TelM="ACTIVE" -acquisition offset failed (acquire 58)

06:08 pointto BS9152 gs=16 PA=180 TelM="ACTIVE" (acquire 59)

06:10 pointto BS9152 gs=13 PA=180 TelM="ACTIVE" (acquire 60)

06:14 pointto BS9152 gs=11 PA=180 TelM="ACTIVE" (acquire 61)

06:17 pointto BS9152 gs=9 PA=180 TelM="ACTIVE" (acquire 62)

06:19 pointto BS9152 gs=5 PA=180 TelM="ACTIVE" (acquire 63)

06:21 pointto BS9152 gs=1 PA=180 TelM="ACTIVE" (acquire 64)

06:27 pointto BS9152 gs=10 PA=0 TelM="ACTIVE" (acquire 65)

06:30 pointto BS9152 gs=8 PA=0 TelM="ACTIVE" (acquire 66)

Fast WFS Data

06:34 pointto M5_0387 gs=0 ACTIVE

IDL collimation through wfscimage 120 0.3 sec exposure

06:42 agw_multiple 50 x 0.3 sec (maybe this is goofy data?) Seeing is 1.3 arcsec at EL=36.

Took images through wfscimage 151.

Close

06:55 Close dome T=-3.9 degC, D=-21 degC, wind (LBTO) ? m/sec.

  • Once GCS is guiding, watch how quickly it determines stable guiding after the acquisition. Previously, this happened after two guide iterations but that was with delays to match the guide rate. I took these delays out which means there is a potential of overcorrecting because of the mount still moving when guide images are being taken. If this is the case it might take longer to ensure stable guiding and in the worst case it might fail during that check. If this is the case, I can quickly add the delays again and ship a patched GCS on Saturday evening that does not have this speed-up attempt in it. (Torsten)

Field Aberration Measurements

BS9122 - Field Aberration Measurements data currently from last night
GStar X Y PA Doug's Angle wfscimage000*
0          
           
           

           

           

           

           

           

           

           




Advanced Planning:

LUCIFER will be available via remote operations in the second half of the nights as needed. If the weather looks bad (but the current forecast is mostly clear) email Nancy by 9:00pm to let her know we won't need LUCIFER support.

Afternoon/Closed-Dome Tests:

Tests that can be done without LUCIFER:

First half of the night:

  • Fast Wavefront Sensing on Bright Star (30min each iteration; Doug, John, Andrew)

It would be useful to gather multiple sets of data in different conditions (equilibrium temps or not, bad seeing vs. good seeing). This test is repeated below.

Second half of the night:

  • Field Aberration Stability Test (2 hr, Andrew)

Persson's Stars Sky Maps

M5 Stars Sky Maps

Star Charts

Persson's Star =
Start time =

Guide Star x y PA Doug's Angle wfscimage000*
- 0 612.5 0 9 1-3
6 0 520.0 43 273 4-6
- 0        
  0        
- 0        
  0        
- 0        
  0        
- 0        
  0        
- 0        
  0        
- 0        
  0        
- 0        
  0        
- 0        
  0        
- 0        
  0        

End time =

  • AGW CCD Orientation recalibration (Jesper, 30m)

Recalibrate the CCD to SFP transformation now that the LUCIFER true north has been determined. This would require a number of exposures of an astrometric field with different values of PA. This would also require 30min of observing time.

  • AGW Non-orthogonality test (Jesper, 30min, details below, requires LUCIFER)

  • Fast Wavefront Sensing on Bright Star (Doug, John, Andrew) - details below

  1. login as LBTO on lbtmu107
  2. make a directory for the data and cd there ("%" is shell prompt)
          % mkdir /Repository/AGW_Data/090116w_fast
          % cd /Repository/AGW_Data/090116w_fast
         
  3. determine the current WFSC integration time
           % getStatus
                  **Note, getStatus reports exposure time in milliseconds
         
  4. Start IDL (">" is the idl prompt)
             % idl
             > @wfsc_init
             > agw_wfsc_multiple, 1, EXP=1       (the first image will probably fail to display: bug in my IDL code)
             > agw_wfsc_multiple, 1, EXP=10     (2nd wfsc image will display)
                    **Note, 1 is number of exposures to take and EXP=10 is integration time in seconds
         
  5. collimate the telescope
          > agw_wfsc_collimate, EXP=30
          > agw_wfsc_collimate, EXP=30
          > ...  As many collimations as is needed to get WFE<400 nm rms
         
  6. if exposure time is good then take at least 60 seconds worth of integration time
          > agw_wfsc_multiple, 6, EXP=10
                ***Note, Crtl-c will stop collection of data and leave everything in happy state
         
  7. repeat step 5-6 above three times
  8. exit IDL
          > exit or Ctrl-d
         
  9. when finished set wfsc exposure time back to original (say 30 seconds)
          % setWFSExpTime 30000
         

Doug's Email Doug's email from Jan. 18 about collecting fast WFSC data

It would be useful to gather multiple sets of data in different conditions (equilibrium temps or not, bad seeing vs. good seeing). This test is repeated above.

  • Collimation model (1-2 hr, low priority, could be skipped)

Tests that need LUCIFER imaging:

  • Re-measure AGW-LUCIFER focus (1 hour, postpone to LUCIFER Com4)
    • Use several different off-axis guide stars are different rotator angles.
    • Use LUCIFER internal focus motion to speed the data acquisition and avoid drifts of temperature or seeing.

Take 15 minutes data in ACTIVE mode in these configurations. If possible, take a series of 5s LUCIFER exposures during each run to compare centroid offsets on LUCIFER to the same data on the guider images. Compare to results of 3a and 3b tests in January.

  • AGW Non-orthogonality test (Jesper, 30min)

We will need a number of LUCIFER images with a probe offset to different positions to confirm that the resulting offsets in the LUCIFER focal plane are indeed linear and not S-shaped.

  • AGW Coordinate Transformation Calibration (Jesper, 2 hours)

Repeat coordinate transform calibrations (2hr): There seems to be a pointing offset in prior data of the order 5mm which I still have not been able to pin-point. It would be very useful to get a new dataset after we have confirmed the rotation axis location in the AGW coordinate system, on LUCIFER, and with a freshly reset telescope pointing.

  • Alignment of chief ray to LUCIFER pupil with M3 (Andrew, Walter, John, 2 hours)
    • Point to bright star
    • Insert LUCIFER pupil viewer
    • Take Pupil image
    • Offset rotator 180 deg
    • Take Pupil image
    • Iterate M3 tip-tilt to minimize the pupil motion
    • Compensate pointing with IE,CA

  • Pointing Model (low priority, 3 hours)

-- JoarBrynnel - 15 Jan 2009
Topic revision: r7 - 03 Feb 2009, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback